# Cooperative Sensing and Communication for ISAC Networks: Performance Analysis and Optimization

### v1

@@ III-A Performance Gain of Cooperative Sensing
$\frac{{2\sum\nolimits_{i=1}^{N}\!{\sum\nolimits_{j=1}^{N}d_{i}^{-\beta}d_{j}^{-\beta}\left(1+\cos\left({{\theta_{i}}-{\theta_{j}}}\right)\right)}}}{{\!\sum\nolimits_{l=1}^{N}{\!\sum\nolimits_{k=1}^{N}{\!\sum\nolimits_{i\geq k}^{N}{\!\sum\nolimits_{j>(k-i)N+l}^{N}\!{(d_{i}d_{j}d_{l}d_{k})^{-\beta}}}}}{{\!\left({{a_{kl}}{b_{ij}}\!-\!{a_{ij}}{b_{kl}}}\right)}^{2}}}}\!\bigg{]}\!,\vspace{-1.5mm}$
where $d_{i}={\left\|{{{\bf{d}}_{i}}}\right\|}^{-\beta}$. To obtain a more tractable CRLB expression, we resort to a simple but tight approximation, and then the following conclusion is proved.
@@
Interestingly, it can be found that the value of expected CRLB in Proposition 1 is only determined by the expected distance from the BS to the typical target. It is verified that (12) achieves a good approximation by Monte Carlo simulations, as shown in Section V. Furthermore, the expected distance from the $n$th closest BS to the typical target can be expressed as $E\left[{{d_{n}}}\right]={{\frac{{\Gamma\left(n+\frac{1}{2}\right)}}{{\lambda_{b}\pi\Gamma(n)}}}}\approx\frac{\sqrt{n}}{\lambda_{b}\pi}$. Then, the CRLB expression can be further approximated as
${\rm{CRLB}}\approx\frac{2}{|\zeta|^{2}{{\lambda_{b}^{2}}{\pi^{2}}\sum\nolimits_{l=1}^{N}{\sum\nolimits_{k=1}^{N}{k^{-\frac{\beta}{2}}l^{-\frac{\beta}{2}}}}}}.\vspace{-1.5mm}$ (13)
@@ VI Conclusion
${\rm{CRLB}}$ $\!=\!{{\rm{E}}_{d}}\!\left[\!{\frac{{2{{\rm{E}}_{\theta}}[{\sum\nolimits_{i=1}^{N}{\sum\nolimits_{j=1}^{N}{{D_{ij}}}(1+\cos({{\theta_{i}}-{\theta_{j}}}))}}]}}{{{{\rm{E}}_{\theta}}[{\sum\nolimits_{l=1}^{N}{\sum\nolimits_{k=1}^{N}{\sum\nolimits_{i\geq k}^{N}{\sum\nolimits_{j>(k-i)N+l}^{N}{{D_{kl}}{D_{ij}}}}}}{Y}}]}}}\!\right]$ (22)

### v2

@@ III-A Performance Gain of Cooperative Sensing
$\frac{{2\sum\nolimits_{i=1}^{N}\!{\sum\nolimits_{j=1}^{N}d_{i}^{-\beta}d_{j}^{-\beta}\left(1+\cos\left({{\theta_{i}}-{\theta_{j}}}\right)\right)}}}{{\!\!\sum\nolimits_{l=1}^{N}{\!\!\sum\nolimits_{k=1}^{N}{\!\sum\nolimits_{i\geq k}^{N}{\!\sum\nolimits^{N}_{j>\!{\lceil(k-i)N+l\rceil}^{+}}\!\!{(d_{i}d_{j}d_{l}d_{k})^{\!-\beta}}}}}{{\!\left({{a_{kl}}{b_{ij}}\!-\!{a_{ij}}{b_{kl}}}\right)}^{2}}}}\!\bigg{]}\!,\vspace{-1.5mm}$
where $d_{i}={\left\|{{{\bf{d}}_{i}}}\right\|}^{-\beta}$ and ${\lceil x\rceil}^{+}=\max(x,1)$. To obtain a more tractable CRLB expression, we resort to a simple but tight approximation, and then the following conclusion is proved.
@@
Interestingly, it can be found that the value of expected CRLB in Proposition 1 is only determined by the expected distance from the BS to the typical target. It is verified that (12) achieves a good approximation by Monte Carlo simulations, as shown in Section V. Furthermore, the expected distance from the $n$th closest BS to the typical target can be expressed as $E\left[{{d_{n}}}\right]={{\frac{{\Gamma\left(n+\frac{1}{2}\right)}}{{\sqrt{\lambda_{b}\pi}\Gamma(n)}}}}\approx\sqrt{\frac{n}{\lambda_{b}\pi}}$. Then, the CRLB expression can be further approximated as
${\rm{CRLB}}\approx\frac{2}{|\zeta|^{2}{{\lambda_{b}^{\beta}}{\pi^{\beta}}\sum\nolimits_{l=1}^{N}{\sum\nolimits_{k=1}^{N}{k^{-\frac{\beta}{2}}l^{-\frac{\beta}{2}}}}}}.\vspace{-1.5mm}$ (13)
@@ VI Conclusion
${\rm{CRLB}}$ $\!=\!{{\rm{E}}_{d}}\!\left[\!{\frac{{2{{\rm{E}}_{\theta}}[{\sum\nolimits_{i=1}^{N}{\sum\nolimits_{j=1}^{N}{{D_{ij}}}(1+\cos({{\theta_{i}}-{\theta_{j}}}))}}]}}{{{{\rm{E}}_{\theta}}[{\sum\nolimits_{l=1}^{N}{\sum\nolimits_{k=1}^{N}{\sum\nolimits_{i\geq k}^{N}{\sum\nolimits_{j>\!{\lceil(k-i)N+l\rceil}^{+}}^{N}\!\!{{D_{kl}}{D_{ij}}}}}}{Y}}]}}}\!\right]$ (22)